[figure: 9.7 cm]

XPS STUDY OF THE ”INFINITE CHAIN” Ca_{2}CuO_{3} AND DISORDER Ca_{0.5}Sr_{0.5}CuO_{2} COMPOUNDS

I. Jacyna-Onyszkiewicz and Yu.B. Poltoracky^{∗} Physics Department, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland ∗NMO ”Monocrystalreactiv”, Lenin 60, 310077 Kharkov, Ukraine

Al Kα excited photoelectron spectra (XPS) of the O 1s, Cu 2p, Ca 2p, Sr 3d core levels have been measured for ”infinite chain” Ca_{2}CuO_{3} and disorder Ca_{0.5}Sr_{0.5}CuO_{2} compounds. For Ca_{2}CuO_{3} the spectrum of the O 1s level was well-fitted with three states at BE’s of 528.9, 531.0 and 532.2 eV respectively, attributable to lattice oxygen in the bulk material (similarly as to HTS oxides), mixed lattice and strongly adsorbed oxygens and OH^{−} species. For Ca_{0.5}Sr_{0.5}CuO_{2} only two state were found for O 1s core level: 529.3 eV and 531.5 eV. For both samples, the Cu 2p 3/2 core level spectra exhibited the same main peak at BE ≈ 934 eV with a prominent shake-up satellite structure, but for Ca_{2}CuO_{3}, an additional component appears which can be connected with copper hydroxide. The valency of the Cu, Ca and Sr atoms is close to +2. The results showed that for the ”infinite chain” Ca_{2}CuO_{3} the surface electronic structures are strongly influenced by a humid atmosphere.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : I. Jacyna-Onyszkiewicz

Address for correspondence : Physics Department, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

Email address : ionys@amu.edu.pl